Calculate the wavelength (in nanometer) associated with a proton moving at 1.0 × 103 m s–1 (Mass of proton = 1.67 × 10–27 kg and h = 6.63 × 10–34 Js) :

Grade:11

2 Answers

Arun
3 years ago

As we learnt in 

De-broglie wavelength -

\lambda = \frac{h}{mv}= \frac{h}{p}

- wherein

where m is the mass of the particle

v its velocity 

p its momentum

 

 \lambda = \frac{h}{mv} = \frac{{6.63} \times {10^{-34}}} {{1.67} \times {10^{-27}} \times {10^3}}

        =0.40 nm

 

Vikas TU
3 years ago
Here,de-broglie wavelength.
wavelenght=h/mv
200× 10^-12=6.626×10^-34/p[1pm=10^-12m,h=6.626×10^-34]
p=6.626×10-34/200×10^-12
p=6.626×10-34/2×10^-10
p=3.313×10^-34+10
p=3.313×10^-24kgms-